OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (amistre64):

3 things to find u dot v the length of u the length of v

OpenStudy (amistre64):

\[cos(\alpha)=\frac{u\cdot v}{|u||v|}\] \[\alpha=\cos^{-1}\left(\frac{u\cdot v}{|u||v|}\right)\]

OpenStudy (anonymous):

u dot v is 46

OpenStudy (anonymous):

i dont know how to do the rest

OpenStudy (amistre64):

the length of a vector is the sqrt of its own dot product |u| = sqrt(u.u) |v| = sqrt(v.v)

OpenStudy (amistre64):

essentially the bottom part is: sqrt(u.u, times v.v)

OpenStudy (anonymous):

so |u| would be sqrt 7?

OpenStudy (amistre64):

sqrt(36+1) = sqrt(37) but yes

OpenStudy (amistre64):

6, -1 6, -1 ----- 36+1 = 37 |u| = sqrt(37)

OpenStudy (anonymous):

ohh so |v| would be sqrt(65)

OpenStudy (amistre64):

correct

OpenStudy (amistre64):

\[\cos(\alpha)=\frac{46}{\sqrt{37*65}}\] then inverse cosine each side

OpenStudy (anonymous):

arccos46/sqrt 37*65 ?

OpenStudy (amistre64):

yes :) which should come out to about 0.353997... radians or 20.28 degrees radians times, 180/pi if you need to convert

OpenStudy (anonymous):

thanks so much!!!!!!!!

OpenStudy (amistre64):

your welcome, and good luck :)